# SelectLLM: Query-Aware Efficient Selection Algorithm for Large Language Models

### v3

@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in solving a wide range of core NLP tasks (Bommasani et al., 2021; Chang et al., 2023). Despite these advances, existing LLMs still struggle with complex tasks such as factually-grounded reasoning and planning (Wei et al., 2022; Kojima et al., 2022; Minaee et al., 2024). Moreover, the wide range of LLMs available seem to exhibit diverse capabilities (Jiang et al., 2023), resulting in no single (especially open-source) LLM being effective across all benchmarks and datasets. Aligning with recent research trends DeepSeek-AI et al. (2025); Mirzadeh et al. (2024), we focus on natural language understanding tasks grounded in reasoning and do not consider natural language generation tasks.
@@ 4.1.2 Data Preparation for the SelectLLM Model: SLData
In this study, we evaluate the performance of each LLM by generating $M$ responses from it for each input query to ensure reliable and replicable behavior of the proposed model. We employ Majority Voting Li et al. (2024) to assess whether a query is correctly answered by the LLM or not. Specifically, Majority Voting (maj@M $\in\{0,1\}$) determines whether the most frequent answer from an LLM matches the gold answer or not. The accuracy with maj@M across all input prompts is reported in Table 1. In the rest of this paper, we consider only those LLMs for which the viable extracted answers are above 90% (see more details in Appendix Section A) to ensure response reliability, resulting in 6 acceptable LLMs for the GSM8K dataset and 7 for the MMLU dataset. We prepare the training dataset for the multi-label classification module of SelectLLM (detailed in the next section) by associating each input query with the LLM(s) whose majority vote answer (across $M$ samples) matches the gold answer, i.e., maj@M $=1$. Formally, the target label for a query prompt $q_{i}\in\mathcal{Q}$ is given by $label\left(q_{i}\right)=\left\{l\>|\>l\in\mathcal{L},maj@M\left(q_{i},l\right)=1\right\}$. We denote this dataset as SLData which is separately prepared for both GSM8K and MMLU.
@@ 4.2 The Proposed SelectLLM Algorithm
    WeightedMaxConf: This policy begins by selecting the top-$s$ LLMs based on their high confidence scores, i.e., for a given query $q_{i}$, we denote the selected LLMs as $L_{s}^{q_{i}}$. Subsequently, we modify the frequency of answer values extracted from the responses of each selected LLM, which involves dividing the frequency of each value by the square root of the confidence score associated with the respective LLM. Finally, we collect all response values and their modified frequencies across the selected LLMs (frequencies are added for the same value). The value with the highest frequency after the modification is selected as the final response. The formal steps are presented in Algorithm 1 in Appendix E. Intuitively, dividing by the square root of the confidence score aims to mitigate biased selection of LLM by the policy similar to Wu et al. (2024). This adjustment ensures fairer opportunities for each selected LLM to contribute to the majority voting.
@@ 6 Results and Discussion
Table 2: Performance of different models on the GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021) test sets. For GSM8K, we use $s=4$, and for MMLU, $s=2$. The value of $s$ is selected based on experiments on the validation set (see Appendix Section I). Acc: with maj@(M $\times$ $L_{s}$) scores reported in percentage (%); Lat: runtime of $M$ generations for a single query (in seconds); MLC: multi-label classifier; $s$: the number of LLMs considered.
@@
Performance of baselines: We evaluated four baseline models – Random, All LLMs, LLM-Blender, and Top-s LLMs – to assess the impact of ensembling LLMs. The performance of the Random baseline model surpasses that of several individual LLMs reported in Fig. 1, demonstrating the potential of utilizing multiple LLMs. Li et al. (2024) reports that increasing the number of LLMs in the model pool generally increases ensemble performance. Interestingly, however, selecting only the top few LLMs in our experiments performs better than including all LLMs (in all LLMs). Thus, the optimal number of LLMs may depend on the diversity of the model pool (i.e., individual capabilities and overall performance on benchmarks). The Top-s LLMs model proves to be the most effective baseline, outperforming even robust approaches such as LLM-Blender, indicating the benefit of prioritizing the best-performing LLMs. However, the latency of most baseline models is high (as they must utilize the entire model pool at inference), limiting their real-world practicality.
@@
The effect of different policies with the SelectLLM algorithm: It can be observed that the LabelledMaxConf policy yields the lowest performance on the GSM8K dataset. This suboptimal performance may be attributed to the policy’s dependence on both the MLC predictions and confidence scores, wherein the classifier predominantly assigns the label metamath-7b-lm. Given that metamath-7b-lm is a specialized model for mathematics and 88% of the SLData training data is associated with this label, this reliance skews the predictions. However, this effect is minimal for the MMLU dataset, where the label distribution across LLMs is more balanced. This limitation is addressed by the MaxConf and WeightedMaxConf, which relax the constraint on MLC label prediction and only operate on confidence scores. This allows models to incorporate other LLMs and push the performance, particularly for the GSM8K data. Mathematical tuning in WeightedMaxConf allows policies to select LLMs more effectively and improve the scores. Overall, the WeightedMaxConf policy emerged as the best performing, with a slight edge over its closest competitor, the MaxConf policy.
@@
SelectLLM vs. baseline models: The proposed SelectLLM model, utilizing the WeightedMaxConf policy, demonstrates superior performance compared to the Random, All LLMs, and LLM-Blender baselines. Additionally, SelectLLM maintains competitive accuracy relative to the Top-$s$ LLMs baseline while achieving significantly lower latency. Specifically, it reduces latency by 13% for the GSM8K dataset and by 70% for the MMLU dataset compared to all baselines (except Random). This substantial efficiency gain underscores the effectiveness of SelectLLM in balancing performance and operational efficiency.
@@
Query awareness analysis of SelectLLM: With this analysis, we aim to understand how the distribution of selected LLMs changes as more LLMs are selected (with increasing values of $s$) in SelectLLM algorithm. Figure 5 in the Appendix presents such distribution. For both datasets, in the top-1 and top-2 subsets, most of the queries are directed to the best-performing LLMs. However, as the subset size increases, the dominance of the top-performing models diminishes, leading to a more uniform distribution where queries are routed towards more LLMs to boost the performance. This indicates the input query awareness of the SelectLLM model, which is adept at assigning a suitable set of LLMs for the input query.
Prediction distribution analysis: Appendix Figure 6 presents the distribution of the number of input queries with correct answer predictions using the top-3 individual LLMs and models with the SelectLLM algorithm across both datasets. It can be observed (count from the right column) that the proposed model is able to correctly provide answers to input queries compared to other individual LLMs, which supports the performance gain reported in Table 2. The distribution also indicates (count from the bottom) that the proposed model utilizes the capabilities of multiple individual LLMs to extract the correct answers. Additionally, Figure 9 illustrates the subject-wise performance differences between SelectLLM and the best performing individual LLM (i.e., gemma-7b-lm) for MMLU. It can be observed that the proposed model shows substantial gains in the majority of subjects, while performing slightly worse for a few of them, showing scope for improvement.
Upper-bound and performance gap with SelectLLM: We measure the maximum performance achievable by SelectLLM with the WeightedMaxConf policy by establishing an upper-bound performance. Specifically, we compute potential subsets based on the labels predicted by the MLC for each query $q_{i}$. A query is considered solved if at least one subset yields a correct answer. The upper bound scores for GSM8K and MMLU are 78.77 and 76.20, respectively. This reveals the following insights: (1) the classifier’s performance is constrained, with weighted F1 scores of 0.71 for GSM8K and 0.68 for MMLU, due to limited training data (approximately 7K for GSM8K and 14K for MMLU), likely contributing to the performance gap between SelectLLM and Oracle. Therefore, augmenting the training data or improving the classifier model could enhance scores. (2) A performance disparity between the best policy and the upper bound suggests the potential for developing better policies (specifically for MMLU). However, since policies rely on the classifier’s confidence scores, enhancing the classifier could also bridge this gap. We will investigate this performance gap from a linguistic perspective in Section 6.1.
Ablation studies: Is the proposed algorithm effective with different LLMs pool sizes (i.e., value of $k$)? To investigate this question, we conduct ablation studies considering various LLM pool sizes, i.e., $k={1,...,6}$, for the GSM8K dataset. We examine two extreme settings: pools with top-$k$ and bottom-$k$ LLMs based on individual LLM performance. This encompasses many configurations as LLMs with similar or different performances may be present in the LLM pool. We also compare this with closet strong Top-$s$ LLM baseline. The results are shown in the Appendix Figure 8. We observe that even with different $k$ values across both top-$k$ and bottom-$k$ setups, the proposed SelectLLM outperforms (in terms of accuracy) the Top-$s$ LLM baseline. This indicates that the proposed approach is LLM pool-agnostic. Moreover, as the number of $k$ values increases, latency becomes a factor: for larger pool sizes, the latency for SelectLLM is much lower than for Top-$s$ LLM. Similar results are observed with the MMLU dataset.
@@ 6.1 Interpretation of Performance Gaps
Quantitative Analysis: We extracted 16 linguistic features specific to input questions, as proposed by Srivatsa and Kochmar (2024) for the GSM8K dataset. In Figure 3, we present the feature value distribution (bar plot) and pair-wise value differences (heatmap) across all four subsets. Major insights are: (1) The feature value distributions for G3 and G4 are notably similar, suggesting that questions only solvable by the Oracle exhibit comparable complexity to those not solved by any LLMs. (2) Features such as length, tree depth, NP count, and argument count emerge as influential across the different subsets, which aligns with the findings of Rabinovich et al. (2023) and Srivatsa and Kochmar (2024). A similar observation holds for the MMLU dataset. Future models should incorporate these features to address the performance gap between SelectLLM and the Oracle.
Qualitative Analysis: We extracted word-level SHAP values [^1] Some samples of interactive SHAP plots are hosted at https://anonymous.4open.science/r/SelectLLM-3621. from the SelectLLM classifier for GSM8K. After mapping key content words to WordNet Miller (1994) synsets, we identified synset groups most detrimental to the solvability of each LLM. Table 3 shows the corresponding SHAP values for GSM8K. See Appendix M for more details. We find that: (1) Not all synset groups affect LLMs uniformly; some are detrimental to one model but beneficial to another (e.g., small numbers & quantifiers). Frequency terms (e.g., ‘‘twice", ‘‘thrice") consistently reduce solvability for most LLMs, but MetaMath remains resilient. (2) Gain using SelectLLM: SelectLLM solves questions with more challenging fractional values and rates and ratios than individual LLMs. (3) Gap between SelectLLM and Oracle: quantifiers, age units, other units, and questions about groups like "family" or "team" relatively increase difficulty in G3 compared to G2. (4) Solved by None: G3/G4 rows show that unsolvable questions contain challenging age units, ordinals (e.g., "first", "third"), and named entities.
@@ 7 Conclusions and Future Directions
It is well known that the one model fits all approach is not universally applicable Huang et al. (2024). For instance, advanced LLMs such as DeepSeek-R1 DeepSeek-AI et al. (2025) demonstrate superior performance in reasoning tasks but exhibit comparatively lower effectiveness in general generation tasks Mercer et al. (2025). Similar observations hold for OpenAI’s and Google’s reasoning models. Building on this observation, we posit that SelectLLM serves as a pivotal step toward the efficient selection of an optimal subset of LLMs from a larger pool, balancing both latency and accuracy, with a focus on reasoning tasks. While the current study does not cover general generation tasks, which are left for future work, the primary challenge for such adaptation lies in implementing robust voting mechanisms. These can be modeled in a simpler manner, as proposed in the All LLMs Li et al. (2023) baseline, to ensure feasibility.
@@ Appendix D Sample Prompts
Appendix E SelectLLM Inference Algorithm
Refer to the Algorithm 1.
[…]
3:for each query $q_{i}$ in $\mathcal{Q}$ do
7:     Initialize an empty set for answer set $A_{i}\leftarrow\phi$
8:     for each $l_{j}$ in $L_{s}$ do
11:         $A_{i}\leftarrow A_{i}\cup a_{i}^{f}$
12:     end for
13:     Return most frequent answer in $A_{i}$
14:end for
@@ Appendix E SelectLLM Inference Algorithm
Appendix F Distribution of Different LLMs in Top-s Subset Selected with SelectLLM

### v4

@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in solving a wide range of core NLP tasks (Bommasani et al., 2021; Chang et al., 2023). Despite these advances, existing LLMs still struggle with complex tasks such as factually-grounded reasoning and planning (Wei et al., 2022; Kojima et al., 2022; Minaee et al., 2024). Moreover, the wide range of LLMs available seem to exhibit diverse capabilities (Jiang et al., 2023), resulting in no single (especially open-source) LLM being effective across all benchmarks and datasets. Aligning with recent research trends DeepSeek-AI et al. (2025); Mirzadeh et al. (2024), we focus on natural language understanding tasks grounded in reasoning, specifically those with discrete output values, and do not consider general natural language generation tasks.
@@ 4.1.2 Data Preparation for the SelectLLM Model: SLData
In this study, we evaluate the performance of each LLM by generating $M$ responses from it for each input query to ensure reliable and replicable behavior of the proposed model. We employ Majority Voting Li et al. (2024) to assess whether a query is correctly answered by the LLM or not. Specifically, Majority Voting(maj@M $\in\{0,1\}$) determines whether the most frequent answer from an LLM matches (using exact string match) the gold answer or not. The accuracy with maj@M across all input prompts is reported in Table 1. In the rest of this paper, we consider only those LLMs for which the viable extracted answers are above 90% (see more details in Appendix Section A) to ensure response reliability, resulting in 6 acceptable LLMs for the GSM8K dataset and 7 for the MMLU dataset. We prepare the training dataset for the multi-label classification module of SelectLLM (detailed in the next section) by associating each input query with the LLM(s) whose majority vote answer (across $M$ samples) matches the gold answer, i.e., maj@M $=1$. Formally, the target label for a query prompt $q_{i}\in\mathcal{Q}$ is given by $label\left(q_{i}\right)=\left\{l\>|\>l\in\mathcal{L},maj@M\left(q_{i},l\right)=1\right\}$. We denote this dataset as SLData which is separately prepared for both GSM8K and MMLU.
[…]
3:for each query $q_{i}$ in $\mathcal{Q}$ do
7:     Initialize an empty set for answer set $A_{i}\leftarrow\phi$
8:     for each $l_{j}$ in $L_{s}$ do
11:         $A_{i}\leftarrow A_{i}\cup a_{i}^{f}$
12:     end for
13:     Return most frequent answer in $A_{i}$
14:end for
@@ 4.2 The Proposed SelectLLM Algorithm
    WeightedMaxConf: This policy begins by selecting the top-$s$ LLMs based on their high confidence scores, i.e., for a given query $q_{i}$, we denote the selected LLMs as $L_{s}^{q_{i}}$. Subsequently, we modify the frequency of answer values extracted from the responses of each selected LLM, which involves dividing the frequency of each value by the square root of the confidence score associated with the respective LLM. Finally, we collect all response values and their modified frequencies across the selected LLMs (frequencies are added for the same value). The value with the highest frequency after the modification is selected as the final response. The formal steps are presented in Algorithm 1. Intuitively, dividing by the square root of the confidence score aims to mitigate biased selection of LLM by the policy similar to Wu et al. (2024). This adjustment ensures fairer opportunities for each selected LLM to contribute to the majority voting.
@@ 6 Results and Discussion
SelectLLM MLC + WeightedMaxConf ($s=1$) 67.24 4.70 63.52 2.97
Table 2: Performance of different models on the GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021) test sets. By default, $s=4$ is used for GSM8K and $s=2$ for MMLU. The value of $s$ is selected based on experiments on the validation set (see Appendix Section 6). Acc: with maj@(M $\times$ $L_{s}$) scores reported in percentage (%); Lat: runtime of $M$ generations for a single query (in seconds); MLC: multi-label classifier; $s$: the number of LLMs considered.
@@
We evaluated four baseline models – Random, All LLMs, LLM-Blender, and Top-s LLMs – to assess the impact of ensembling LLMs. The performance of the Random baseline model surpasses that of several individual LLMs reported in Fig. 1, demonstrating the potential of utilizing multiple LLMs. Li et al. (2024) reports that increasing the number of LLMs in the model pool generally increases ensemble performance. Interestingly, however, selecting only the top few LLMs in our experiments performs better than including all LLMs (in all LLMs). Thus, the optimal number of LLMs may depend on the diversity of the model pool (i.e., individual capabilities and overall performance on benchmarks). The Top-s LLMs model proves to be the most effective baseline, outperforming even robust approaches such as LLM-Blender, indicating the benefit of prioritizing the best-performing LLMs. However, the latency of most baseline models is high (as they must utilize the entire model pool at inference), limiting their real-world practicality.
The effect of different policies with the
@@
SelectLLM algorithm: It can be observed that the LabelledMaxConf policy yields the lowest performance on the GSM8K dataset. This suboptimal performance may be attributed to the policy’s dependence on both the MLC predictions and confidence scores, wherein the classifier predominantly assigns the label metamath-7b-lm. Given that metamath-7b-lm is a specialized model for mathematics and 88% of the SLData training data is associated with this label, this reliance skews the predictions. However, this effect is minimal for the MMLU dataset, where the label distribution across LLMs is more balanced. This limitation is addressed by the MaxConf and WeightedMaxConf, which relax the constraint on MLC label prediction and only operate on confidence scores. This allows models to incorporate other LLMs and push the performance, particularly for the GSM8K data. Mathematical tuning in WeightedMaxConf allows policies to select LLMs more effectively and improve the scores. Overall, the WeightedMaxConf policy emerged as the best performing, with a slight edge over its closest competitor, the MaxConf policy.
SelectLLM
@@
vs. baseline models: The proposed SelectLLM model, utilizing the WeightedMaxConf policy, demonstrates superior performance compared to the Random, All LLMs, and LLM-Blender baselines. Additionally, SelectLLM maintains competitive accuracy relative to the Top-$s$ LLMs baseline while achieving significantly lower latency. Specifically, it reduces latency by 13% for the GSM8K dataset and by 70% for the MMLU dataset compared to all baselines (except Random). This substantial efficiency gain underscores the effectiveness of SelectLLM in balancing performance and operational efficiency.
Query awareness analysis of SelectLLM:
With this analysis, we aim to understand how the distribution of selected LLMs changes as more LLMs are selected (with increasing values of $s$) in SelectLLM algorithm. Figure 6 in the Appendix presents such distribution. For both datasets, in the top-1 and top-2 subsets, most of the queries are directed to the best-performing LLMs. However, as the subset size increases, the dominance of the top-performing models diminishes, leading to a more uniform distribution where queries are routed towards more LLMs to boost the performance. This indicates the input query awareness of the SelectLLM model, which is adept at assigning a suitable set of LLMs for the input query.
Prediction distribution analysis:
Appendix Figure 7 presents the distribution of the number of input queries with correct answer predictions using the top-3 individual LLMs and models with the SelectLLM algorithm across both datasets. It can be observed (count from the right column) that the proposed model is able to correctly provide answers to input queries compared to other individual LLMs, which supports the performance gain reported in Table 2. The distribution also indicates (count from the bottom) that the proposed model utilizes the capabilities of multiple individual LLMs to extract the correct answers. Additionally, Figure 9 illustrates the subject-wise performance differences between SelectLLM and the best performing individual LLM (i.e., gemma-7b-lm) for MMLU. It can be observed that the proposed model shows substantial gains in the majority of subjects, while performing slightly worse for a few of them, showing scope for improvement.
Upper-bound and performance gap with SelectLLM:
We measure the maximum performance achievable by SelectLLM with the WeightedMaxConf policy by establishing an upper-bound performance. Specifically, we compute potential subsets based on the labels predicted by the MLC for each query $q_{i}$. A query is considered solved if at least one subset yields a correct answer. The upper bound scores for GSM8K and MMLU are 78.77 and 76.20, respectively. This reveals the following insights: (1) the classifier’s performance is constrained, with weighted F1 scores of 0.71 for GSM8K and 0.68 for MMLU, due to limited training data (approximately 7K for GSM8K and 14K for MMLU), likely contributing to the performance gap between SelectLLM and Oracle. Therefore, augmenting the training data or improving the classifier model could enhance scores. (2) A performance disparity between the best policy and the upper bound suggests the potential for developing better policies (specifically for MMLU). However, since policies rely on the classifier’s confidence scores, enhancing the classifier could also bridge this gap. We will investigate this performance gap from a linguistic perspective in Section 6.1.
Ablation studies:
Is the proposed algorithm effective with different LLMs pool sizes (i.e., value of $k$)? To investigate this question, we conduct ablation studies considering various LLM pool sizes, i.e., $k={1,...,6}$, for the GSM8K dataset. We examine two extreme settings: pools with top-$k$ and bottom-$k$ LLMs based on individual LLM performance. This encompasses many configurations as LLMs with similar or different performances may be present in the LLM pool. We also compare this with closet strong Top-$s$ LLM baseline. The results are shown in the Appendix Figure 8. We observe that even with different $k$ values across both top-$k$ and bottom-$k$ setups, the proposed SelectLLM outperforms (in terms of accuracy) the Top-$s$ LLM baseline. This indicates that the proposed approach is LLM pool-agnostic. Moreover, as the number of $k$ values increases, latency becomes a factor: for larger pool sizes, the latency for SelectLLM is much lower than for Top-$s$ LLM. Similar results are observed with the MMLU dataset.
@@
Latency vs. Accuracy vs. $s$-Value:
Figure 4 shows the relationship between latency, accuracy, and $s$-values for different SelectLLM policies across both datasets along with the corresponding development sets. WeightedMaxConf consistently outperforms both MaxConf and Random in terms of accuracy and latency for most $s$-values (except for 5 and 6 on GSM8K). This highlights the superiority of WeightedMaxConf, maintaining its effectiveness even with a small number of LLMs. A larger number of LLMs and higher latency are required for GSM8K, whereas a lower $s$-value and latency suffice for MMLU to achieve high accuracy. Selecting 4 LLMs for GSM8K and 2 for MMLU yields optimal performance, demonstrating the effectiveness and efficiency of SelectLLM with WeightedMaxConf. Using this optimal value of $s$ results on the test set are reported in Table 2.
Additionally, we also report the performance of SelectLLM with WeightedMaxConf for $s=1$ in Table 2 to understand the impact of selecting the best query-aware LLM. The results indicate that there is a large performance gap when selecting multiple LLMs, underscoring that ensembling is a promising direction.
@@
Figure 4: Latency vs. accuracy vs. $s$-values plots for different SelectLLM policies on GSM8K (left) and MMLU (right) development sets. $s$-value: total number of LLMs selected with SelectLLM model.
@@ 6.1 Interpretation of Performance Gaps
Quantitative Analysis:
We extracted 16 linguistic features specific to input questions, as proposed by Srivatsa and Kochmar (2024) for the GSM8K dataset. In Figure 3, we present the feature value distribution (bar plot) and pair-wise value differences (heatmap) across all four subsets. Major insights are: (1) The feature value distributions for G3 and G4 are notably similar, suggesting that questions only solvable by the Oracle exhibit comparable complexity to those not solved by any LLMs. (2) Features such as length, tree depth, NP count, and argument count emerge as influential across the different subsets, which aligns with the findings of Rabinovich et al. (2023) and Srivatsa and Kochmar (2024). A similar observation holds for the MMLU dataset. Future models should incorporate these features to address the performance gap between SelectLLM and the Oracle.
Qualitative Analysis:
We extracted word-level SHAP values [^1] Some samples of interactive SHAP plots are hosted at https://anonymous.4open.science/r/SelectLLM-3621. from the SelectLLM classifier for GSM8K. After mapping key content words to WordNet Miller (1994) synsets, we identified synset groups most detrimental to the solvability of each LLM. Table 3 shows the corresponding SHAP values for GSM8K. See Appendix K for more details. We find that: (1) Not all synset groups affect LLMs uniformly; some are detrimental to one model but beneficial to another (e.g., small numbers & quantifiers). Frequency terms (e.g., ‘‘twice", ‘‘thrice") consistently reduce solvability for most LLMs, but MetaMath remains resilient. (2) Gain using SelectLLM: SelectLLM solves questions with more challenging fractional values and rates and ratios than individual LLMs. (3) Gap between SelectLLM and Oracle: quantifiers, age units, other units, and questions about groups like "family" or "team" relatively increase difficulty in G3 compared to G2. (4) Solved by None: G3/G4 rows show that unsolvable questions contain challenging age units, ordinals (e.g., "first", "third"), and named entities.
@@ 7 Conclusions and Future Directions
It is well known that the one model fits all approach is not universally applicable Huang et al. (2024). For instance, advanced LLMs such as DeepSeek-R1 DeepSeek-AI et al. (2025) demonstrate superior performance in reasoning tasks but exhibit comparatively lower effectiveness in general generation tasks Mercer et al. (2025). Similar observations hold for OpenAI’s and Google’s reasoning models. Building on this observation, we posit that SelectLLM serves as a pivotal step toward the efficient selection of an optimal subset of LLMs from a larger pool, balancing both latency and accuracy, with a focus on reasoning tasks. While the current study does not cover general generation tasks, which are left for future work, the primary challenge for such adaptation lies in implementing robust voting mechanisms. These can be modeled in a simpler manner, as proposed in the All LLMs Li et al. (2023) baseline, to ensure feasibility. Similarly, the idea proposed by Odumakinde et al. (2024) could be adapted for general generation tasks.
@@ Appendix D Sample Prompts
Appendix E Distribution of Different LLMs in Top-$s$ Subset Selected with SelectLLM